# Connecting the Dots: Is Mode-Connectedness the Key to Feasible Sample-Based Inference in Bayesian Neural Networks?

### v1

@@ 5 Multimodality of Posteriors
More movement deeper down. It turns out, however, that this phenomenon requires a more nuanced analysis. As shown in Fig. 2, the posterior density is heavily concentrated on disconnected modes in the first layer, but then becomes more diffuse and thus connected in deeper parts of the BNN, before concentrating again in the output layer. Our observation is reflected by chains becoming notably more variable in deeper layers (cf. Fig. 3 and 4); only in the last layer does the variance decrease again. The explanation for this behavior can be found in the varying degree of overparameterization throughout the network. We conjecture that disconnected modes occur mainly in layers with limited flexibility, i.e., where the role of the neurons is controlled by their close connection to the input (or output) that populates a certain manifold. Hidden-layer weights, by contrast, have vastly more degrees of freedom in assuming diverse values and swapping their roles in creating input embeddings as long as the output is realized in an appropriate range.
The layerwise investigation has interesting relations to the mode connectivity studies in Izmailov et al. (2021). We believe our results could shed light on the effectiveness of methods such as subspace inference (Izmailov et al., 2020; Dold et al., 2024) that move or sample in directions of quasi-constant loss. This hypothesis needs further evaluation. More immediately, as we discuss in the next section, our observations have important consequences for practical SBI.
@@ 6.1 How to Handle Multimodality in Practice?
Insight: Multiple chains with multiple samples improve predictive performance and uncertainty quantification.
Performance and uncertainty. Fig. 5 (further examples in Appendix C.3) shows a clear upward trend of LPPD in both the number of chains and the number of samples. The same general pattern can be observed for the RMSE (Appendix C.3). Notably, the LPPD increases more strongly for multiple chains than the RMSE, suggesting the importance of covering multiple modes, but we also see performance gains by exploring the region around modes with more samples. It seems promising that we obtain good results with a maximum of 12 chains despite the large number of symmetries. This is in line with findings from Wiese et al. (2023) who ran more than 1,200 chains, but also observed early saturation of performance metrics.
Predictive coverage. Bayesian inference enables the computation of predictive credibility intervals, which should be calibrated and cover the true prediction. Fig. 6 (further results in Appendix C.3) shows the coverage of credibility intervals for varying amounts of samples (left) and chains (right). For a calibrated model (represented by the diagonal), the observed labels fall within the $\alpha\%$ credibility interval in $\alpha\%$ of cases. With increasing chains and samples, we can see a trend from overconfidence to close-to-nominal coverage. These results confirm that the use of multiple chains and samples indeed produces well-calibrated posteriors.
@@ 6.2 Convergence Diagnostics
Parameter space convergence. The classical $\widehat{R}$ diagnostic assumes identifiable parameters, which is not given in BNNs due to the aforementioned symmetries. Our experiments in App. B also confirm empirically that $\widehat{R}$ is not a suitable metric. Another pitfall we can infer from our results is the aggregation of $\widehat{R}$ over all model parameters. As $\widehat{R}$ is normalized using the within-chain variance, it decreases in the later layers of the BNN due to the increase in variance (Fig. 4), leading to the false conclusion of convergence of the whole network when averaging across all weights.
@@
Therefore, we propose to monitor the convergence of each chain individually by the cumulative LPPD as defined in Eq. 8 in the Appendix and shown in, e.g., Fig. 7 (right). By running multiple chains, we can compare the different cumulative LPPD paths and thereby obtain a better understanding of the convergence of each chain and the difference in chain performance. This can also be used as an early-stopping criterion of the sampling process for potentially converged or diverging chains, freeing resources and allowing to start new chains for more efficiency.
@@
7 Dying Sampler & Bayesian Deep Ensembles
@@ 7.1 The Dying Sampler Problem
Although Section 4 established that the use of bounded activation functions largely avoids the dying sampler problem, it would be unsatisfactory to simply dismiss ReLU BNNs as infeasible or refrain from using larger architectures. While previously discussed results on prior variance offer a possible explanation for this phenomenon, a—if not the most important—cause is already evident in Fig. 1: Starting values for chains might force the sampler to get stuck or not explore meaningful hypotheses.
For example, as a consequence of the ReLU activation function being zero for negative input signals, many neurons, just by chance, will not be activated or confined to regions of very small probability (the white area in Fig. 1). Subsequent proposals will then also end up in regions with a posterior probability of practically zero. We can thus draw an analogy to the “dying ReLU problem” (Lu et al., 2020) caused by zero gradients of non-activated ReLU neurons in optimization.
7.2 Bayesian Deep Ensembles
As already elaborated in Section 4, using a different initialization solves convergence issues such as the dying sampler problem. This motivates our concluding proposal for practically useful SBI, which we call Bayesian deep ensembles (BDEs) inspired by their similarity to DEs. Specifically, we suggest running a standard optimization procedure for $M$ networks in a non-Bayesian fashion, and using the resulting $M$ sets of network weights as initial proposals for the samplers. Note that this does not require changing the prior distribution. We combine the information of all chains in the classical Bayesian sense by merging their empirical distributions. While this will coincide with DEs when using a one-sample BDE and estimating the posterior mean, other posterior characteristics generally differ. BDEs have several attractive properties, which we discuss in the following.
Table 3: Average RMSE and LPPD values ($\pm$ their standard deviations) of the LM, a classic DNN, a DE, and a BDE (i.e., a BNN with DNN warm starts) using different amounts of samples (in brackets) for all benchmark data sets (in rows; abbreviated by their first letter). All networks have two hidden layers with 16 neurons each and ReLU activations. The best method per data set is highlighted in bold.
@@ 7.2 Bayesian Deep Ensembles
LM DNN DE BDE (10) BDE (100) BDE (1000) LM DNN DE BDE (10) BDE (100) BDE (1000)
@@
Expected improvement. BDEs are equivalent to DEs in the edge case where we only obtain one sample from the posterior (the initial weight). As DEs have repeatedly shown top-notch quantifying quantification, this means that BDEs are likely at least as good as DEs in this respect. In our experiments, BDEs indeed provide better LPPD for as few as ten samples (cf. Table 3).
A modular Bayesian toolbox. Similar to the Laplace approximation (Daxberger et al., 2021), BDEs can be used post-optimization, allowing to combine prior knowledge of any kind with a non-Bayesian network into a BNN. In particular, BDEs are applicable to pre-trained networks.
Flexibility. While BDEs are as expensive as DEs per design and require further computation for additional sampling, this second process can be flexibly adjusted to the availability of resources and computing time. Furthermore, since our previous results suggest that a few samples already yield very good performance when using multiple chains, it might be possible to shorten the sampling step considerably by saving on warmup iterations.
@@ 7.3 Numerical Experiments
In order to investigate the effectiveness of BDEs, we run an additional benchmark study. Proving that the proposed warm start can solve the dying sampler problem, we use BNNs with ReLU activation and compare the performance of BDEs using all chains to those of an equivalent DNN and DE. In other words, we do not need to filter chains for the comparisons as done in Section 4.
Results. We summarize our findings in Table 3. As becomes evident from the comparisons with the LM and DE, the warm starts of samplers using the $M$ different network solutions avoid any dying sampler problems (no model shows worse performance than the baseline LM). Furthermore, we see the added benefit of sampling around the DE solutions when continuing the sampling from these solutions after a very short warmup phase of 100 steps, suggesting that this process supplies valuable information to the model both for predictive (RMSE) and uncertainty (LPPD) performance. While the best cold-started BNNs required 10,000 warmup steps, the BDEs achieve competitive performance with a 100$\times$ shorter warmup phase. We thus conclude that BDEs, though conceptually simple, offer consistent and robust improvement over classical BNNs.
@@ 8 Discussion
In this work, we discussed sampling-based inference in BNNs. We argue that sampling from a posterior in these networks is not only feasible but can achieve SOTA results when accounting for the nature of posterior landscapes. Through extensive experiments with different BNN architectures, we present insights into the successes and limitations of SBI in this context. Critical findings include the performance differences of samplers for bounded and unbounded activation functions and the increasing connectedness of modes in deeper layers. Following these results, we recommend a multi-chain and multi-sample strategy using NUTS, with a convergence diagnostic that accounts for heterogeneous variances in different layers. In order to allow samplers to also explore highly multimodal landscapes, particularly those induced by ReLU networks, we propose a novel approach, Bayesian deep ensembles (BDEs), which use an optimized network as a starting point for posterior sampling. These ensembles can be seen as a modular Bayesian toolbox applicable to any network post-optimization independent of the choice of prior knowledge.
Limitations and future work. Due to the multitude of experiments, as well as the long duration of sampling procedures, the analysis presented in this paper is limited in some aspects. In particular, only full-batch sampling routines were tested. An exciting future analysis is therefore to explore the potential of SG-MCMC samplers in the context of BDEs, as well as an extension to larger data sets. Likewise, the performance of BDEs in uncertainty-related downstream tasks, such as out-of-distribution detection, remains an open question we plan to address in the future. Finally, it would be interesting to deepen our findings in conjunction with the insights from mode connectivity (Garipov et al., 2018) and subspace research (Izmailov et al., 2020).
Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ B.2 Function Space
Figure 10: Histograms, mean and standard deviation of the parameter-space (PS) $\widehat{R}^{(2)}$ and function-space (FS) $\widehat{R}^{(2)}_{\psi_{\text{PSC}}}$, evaluated for each parameter and test data entry respectively. For each data set, all fitted chains with better than LM RMSE performance are used. In all cases, the BNN consists of two hidden layers with 16 neurons each and tanh activation. NUTS with a warmup phase of 10k steps and 8k samples per chain as well as unit Gaussian priors are used.
@@
Expanding-window function-space convergence. Convergence diagnostics usually evaluate chain mixing post-hoc. We propose an expanding-window approach, enabled by chain-wise analysis, that allows us to assess chain-wise convergence and between-chain agreement in function space at any time point during the sampling process. In particular, this entails the option of online resource management by early stopping of chains: when the sampler shows evidence of being stuck and risks wasting compute, the chain can be restarted.
@@ C.1 General Feasibility
Table 4 contains the results reported in Table 1 with additional standard deviations in parentheses. The architecture is a two-hidden layer MLP with 16 neurons in each hidden layer. The BNNs are sampled using 10,000 warmup steps of NUTS and 8,000 samples for each of the 12 chains. Only chains that proved to be better than the baseline LM are considered for the calculation (here always at least 10). The tanh activation function and unit Gaussian priors are applied. Experiments reported in Table 2 also use unit Gaussian priors.
@@
data set LM RF DNN DE BNN (RS) BNN
airfoil 0.716 (0.022) 0.288 (0.005) 0.252 (0.024) 0.239 (0.023) 0.25 (0.009) 0.182 (0.009)
bikesharing 0.790 (0.028) 0.304 (0.031) 0.374 (0.002) 0.365 (0.003) 0.362 (0.013) 0.253 (0.036)
concrete 0.630 (0.018) 0.312 (0.018) 0.317 (0.015) 0.282 (0.018) 0.554 (0.158) 0.258 (0.020)
@@ C.2 Mulimodality of Posteriors
The traceplots displayed in Figures 17 and 18 are both examples of the fact that symmetry-induced multimodality exists and also illustrate how the within-chain variance increases in deeper layers. Moreover, Figure 18 shows how for the same setting some (in this case two) chains are “dying”, i.e., have a near-constant trace plot, while others explore the space in a meaningful way (implied by the good performance of the induced BNN).

### v2

@@ 5 Multimodality of Posteriors
More movement deeper down. It turns out, however, that this phenomenon requires a more nuanced analysis. As shown in Fig. 2, the posterior density is heavily concentrated on disconnected modes in the first layer, but then becomes more diffuse and thus connected in deeper parts of the BNN before concentrating again in the output layer. Our observation is reflected by chains growing notably more variable in deeper layers (cf. Fig. 3 and 4); only in the last layer does the variance decrease again. The explanation for this behavior can be found in the varying degree of overparameterization throughout the network. We conjecture that disconnected modes occur mainly in layers with limited flexibility, i.e., where the role of the neurons is controlled by their close connection to the input (or output) that populates a certain manifold. Hidden-layer weights, by contrast, have vastly more degrees of freedom in assuming diverse values and swapping their roles in creating input embeddings as long as the output is realized in an appropriate range.
Role of network layers. Our intuition about the roles of different layers echoes recent findings about the varying robustness of layers to re-initialization during training (where the initial layer, in accordance with our results, proves much more sensitive to resetting the parameter values mid-training than intermediate ones; Zhang et al., 2022). The layerwise investigation also has interesting relations to the mode connectivity studies in Izmailov et al. (2021). Taking the notion of mode connectivity to a per-layer level, work concurrent to ours (Adilova et al., 2024) suggests that intermediate layers play a different role for the success of layerwise neuron alignment (interpolating between multiple independent solutions; cf. Ainsworth et al., 2023) than the first or last one. We further believe our results could shed light on the effectiveness of methods such as subspace inference (Izmailov et al., 2020; Dold et al., 2024) that move or sample in directions of quasi-constant loss. More immediately, as we discuss in the next section, they have important consequences for practical SBI.
@@ 6.1 How to Handle Multimodality in Practice?
Insight: Multiple samples from multiple chains improve predictive performance and uncertainty quantification.
Performance and uncertainty. Fig. 5 (further examples in Section C.3) shows a clear upward trend of LPPD in both the number of chains and the number of samples. The same general pattern can be observed for the RMSE (Appendix C.3). Notably, the LPPD increases more strongly for multiple chains than the RMSE, suggesting the importance of covering multiple modes, but we also see performance gains by exploring the region around modes with more samples. It seems promising that we obtain good results with a maximum of 12 chains despite the large number of symmetries. This is in line with findings from Wiese et al. (2023) who ran more than 1,200 chains but also observed early saturation of performance metrics.
Predictive coverage. Bayesian inference enables the computation of predictive credibility intervals which should be calibrated and cover the true prediction. Fig. 6 (further results in Section C.3) shows the coverage of credibility intervals for varying amounts of samples (left) and chains (right). For a calibrated model (represented by the diagonal), the observed labels fall within the $\alpha\%$ credibility interval in $\alpha\%$ of cases. With increasing chains and samples, we can see a trend from overconfidence to close-to-nominal coverage. These results confirm that the use of multiple chains and samples indeed produces well-calibrated posteriors.
@@ 6.2 Convergence Diagnostics
Parameter space convergence. The classical $\widehat{R}$ diagnostic assumes identifiable parameters, which is not given in BNNs due to the aforementioned symmetries. Our experiments in Appendix B also confirm empirically that $\widehat{R}$ is not a suitable metric. Another pitfall we can infer from our results is the aggregation of $\widehat{R}$ over all model parameters. As $\widehat{R}$ is normalized using the within-chain variance, it decreases in the later layers of the BNN due to the increase in variance (Fig. 4), leading to the false conclusion of convergence of the whole network when averaging across all weights.
@@
Therefore, we propose to monitor the convergence of each chain individually by the cumulative LPPD, as defined in Eq. 8 in the Appendix and shown in, e.g., Fig. 7 (right). By running multiple chains, we can compare the different cumulative LPPD paths and thereby obtain a better understanding of the convergence of each chain and the difference in chain performance. This can also be used as an early-stopping criterion of the sampling process for potentially converged or diverging chains, freeing resources and allowing to start new chains for more efficiency.
@@
7 Dying Sampler & Deep Ensemble Initialization
@@ 7.1 The Dying Sampler Problem
Although Section 4 established that the use of bounded activation functions largely avoids the dying sampler problem, it would be unsatisfactory to simply dismiss ReLU BNNs as infeasible or refrain from using larger architectures. While previously discussed results on prior variance offer a possible explanation for this phenomenon, a—if not the most important—cause is already evident in Fig. 1: Starting values for chains might prompt the sampler to get stuck or not explore meaningful hypotheses.
For example, as a consequence of the ReLU activation function being zero for negative input signals, many neurons—just by chance—will not be activated, creating low-density regions at the locations of the weights associated with those neurons (the white area in Fig. 1). Subsequent proposals will then also end up in regions with a posterior probability of practically zero. We can thus draw an analogy to the “dying ReLU problem” (Lu et al., 2020) caused by zero gradients of non-activated ReLU neurons in optimization.
7.2 Deep Ensemble Initialized MCMC
Table 3: Average RMSE and LPPD values ($\pm$ their standard deviations) of the LM, a classic DNN, a DE, and our method (i.e., DEI-MCMC) using different amounts of samples (in brackets) for all benchmark data sets (in rows; abbreviated by their first letter). All networks have two hidden layers with 16 neurons each and ReLU activations. The best method per data set is highlighted in bold.
@@ 7.2 Deep Ensemble Initialized MCMC
LM DNN DE Ours (10) Ours (100) Ours (1000) LM DNN DE Ours (10) Ours (100) Ours (1000)
@@
As already elaborated in Section 4, using a different initialization solves convergence issues such as the dying sampler problem. This motivates our concluding proposal for practically useful SBI, which we call Deep Ensemble Initialized MCMC (DEI-MCMC). Inspired by their similarity to DEs one could also think of them as a Bayesian Deep Ensemble. Specifically, we suggest running a standard optimization procedure for $M$ networks in a non-Bayesian fashion, and using the resulting $M$ sets of network weights as initial proposals for the samplers. Note that this does not require changing the prior distribution. We combine the information of all chains in the classical Bayesian sense by merging their empirical distributions. DEI-MCMC has several attractive properties, which we discuss in the following.
@@
Expected improvement. DEI-MCMC is equivalent to DEs in the edge case where we only obtain one sample from the posterior (the initial weight). As DEs have repeatedly shown top-notch quantifying quantification, this means that DEI-MCMC is likely at least as good as DEs in this respect. In our experiments, DEI-MCMC indeed provides better LPPD for as few as ten samples (cf. Table 3).
A modular Bayesian toolbox. Similar to the Laplace approximation (Daxberger et al., 2021), DEI-MCMC can be used post-optimization, allowing to combine prior knowledge of any kind with a non-Bayesian network into a BNN. In particular, DEI-MCMC is applicable to pre-trained networks.
Flexibility. While DEI-MCMC is as expensive as DEs by design and requires further computation for additional sampling, this second process can be flexibly adjusted to the availability of resources and computing time. Furthermore, since our previous results suggest that a few samples already yield very good performance when using multiple chains, it might be possible to shorten the sampling step considerably by saving on warmup iterations.
@@ 7.3 Numerical Experiments
In order to investigate the effectiveness of DEI-MCMC, we run an additional benchmark study. Proving that the proposed warm start can solve the dying sampler problem, we use BNNs with ReLU activation and compare the performance of DEI-MCMC using all chains to those of an equivalent DNN and DE. In other words, we do not need to filter chains for the comparisons as done in Section 4.
Results. We summarize our findings in Table 3. As becomes evident from the comparisons with the LM and DE, the warm starts of samplers using the $M$ different network solutions avoid any dying sampler problems (no model shows worse performance than the baseline LM). Furthermore, we see the added benefit of sampling around the DE solutions when continuing the sampling from these solutions after a very short warmup phase of 100 steps, suggesting that this process supplies valuable information to the model both for predictive (RMSE) and uncertainty (LPPD) performance. It should be noted that, although DEI-MCMC tends to benefit from longer sampling, there is a diminishing return to extending the sampling phase beyond the 1,000 samples reported in the last column of Table 3, with only marginal improvements from drawing even up to 10,000 samples. While the best cold-started BNNs required 10,000 warmup steps, DEI-MCMC achieves competitive performance with a 100$\times$ shorter warmup phase. We thus conclude that DEI-MCMC, though conceptually simple, offers a consistent and robust improvement over classical BNNs.
@@ 8 Discussion
In this work, we discussed sampling-based inference in BNNs. We argue that sampling from a posterior in these networks is not only feasible but can achieve SOTA results when accounting for the nature of posterior landscapes. Through extensive experiments with different BNN architectures, we present insights into the successes and limitations of SBI in this context. Critical findings include performance differences of samplers for bounded and unbounded activation functions and the increasing connectedness of modes in deeper layers. Following these results, we recommend a multi-chain and multi-sample strategy using NUTS, with a convergence diagnostic that accounts for heterogeneous variances in different layers. In order to allow samplers to also explore highly multimodal landscapes, particularly those induced by ReLU networks, we propose a novel approach, Deep Ensemble Initialized MCMC (DEI-MCMC), which uses optimized networks as a starting point for posterior sampling. These ensembles can be seen as a modular Bayesian toolbox applicable to any network post-optimization independent of the state of prior knowledge.
Limitations and future work. Due to the multitude of experiments, as well as the long duration of sampling procedures, the analysis presented in this paper is limited in some aspects. In particular, only full-batch sampling routines were tested. An exciting target for future analysis, therefore, is to explore the potential of SG-MCMC samplers in the context of DEI-MCMC, as well as an extension to larger data sets. Likewise, the performance of DEI-MCMC in uncertainty-related downstream tasks, such as out-of-distribution detection, remains an open question. It would further be interesting to deepen our findings in conjunction with the insights from mode connectivity (Garipov et al., 2018) and subspace research (Izmailov et al., 2020). Finally, inspired by our tentative results and the work by Adilova et al. (2024), all these avenues of future work should be pursued adopting a layerwise lens.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ B.2 Function Space
Figure 10: Histograms, mean and standard deviation of the parameter space (PS) $\widehat{R}^{(2)}$ and function space (FS) $\widehat{R}^{(2)}_{\psi_{\text{PSC}}}$, evaluated for each parameter and test data entry respectively. For each data set, all fitted chains with better than LM RMSE performance are used. In all cases, the BNN consists of two hidden layers with 16 neurons each and tanh activation. NUTS with a warmup phase of 10k steps and 8k samples per chain as well as unit Gaussian priors are used.
@@
Expanding-window function space convergence. Convergence diagnostics usually evaluate chain mixing post-hoc. We propose an expanding-window approach, enabled by chain-wise analysis, that allows us to assess chain-wise convergence and between-chain agreement in function space at any time point during the sampling process. In particular, this entails the option of online resource management by early stopping of chains: when the sampler shows evidence of being stuck and risks wasting compute, the chain can be restarted.
@@ C.1 General Feasibility
Table 4 contains the results reported in Table 1 with additional standard deviations in parentheses. The architecture is a two-hidden layer MLP with 16 neurons in each hidden layer. The BNNs are sampled using 10,000 warmup steps of NUTS and 8,000 samples for each of the 12 chains. Only chains that proved to be better than the baseline LM are considered for the calculation (here always at least 10). The tanh activation function and unit Gaussian priors are applied. Experiments reported in Table 2 also use unit Gaussian priors. Note that for the Random Forest baseline an extensive HPO search using optuna (Akiba et al., 2019) is performed. Thus, the comparison is rather conservative as we do not perform a neural architecture search for the network-based models. The relevant yardstick in this comparison are the Deep Ensembles which share the same inductive biases as the BNNs.
@@
data set LM RF (tuned) DNN DE BNN (RS) BNN
airfoil 0.716 (0.022) 0.255 (0.005) 0.252 (0.024) 0.239 (0.023) 0.25 (0.009) 0.182 (0.009)
bikesharing 0.790 (0.028) 0.231 (0.007) 0.374 (0.002) 0.365 (0.003) 0.362 (0.013) 0.253 (0.036)
concrete 0.630 (0.018) 0.304 (0.021) 0.317 (0.015) 0.282 (0.018) 0.554 (0.158) 0.258 (0.020)
@@ C.2 Mulimodality of Posteriors
In addition to the analyses reported in Figures 2, 3 and 4, we investigate the sampling path for each chain. Here, instead of aggregating insights from marginal parameter traces, we take a joint view on the sampling path. In order reduce the dimensionality of this problem, we perform a principal component analysis (PCA; Wold et al., 1987) given $S$ posterior samples from the $k$-th chain $\bm{\Theta}=(\bm{\theta}^{(k,1)},\bm{\theta}^{(k,2)},\dots,\bm{\theta}^{(k,S)})^{\top}$ Taking the three first principal components (each of dimension $|\bm{\theta}^{(k,s)}|$), we can project the sampling path into three dimensions. The first three principal components explain a substantial amount of variance of the sampling path ($>50\%$), which suggests that insights from the PCA-approximated path are indicative of the original, high-dimensional situation. This is confirmed visually, e.g., in Fig. 17, which is a representative example for the first chain of the energy dataset. Summing up the element-wise absolute values of the orthonormal vectors yields a proxy of how important the particular parameter is for the movement of the sampler. Applying this methodology and grouping the PCA factor loadings layer-wise, as well as averaging across experiments and chains, we obtain Fig. 18. The same pattern as in Fig. 4 emerges, i.e., deeper layers are more important for the movement of the chain in the parameter space. We conclude from this that our findings pertain to a robust pattern.
Figure 17: Example sampling path of the first chain of the energy dataset projected in three dimensions using PCA. Sampler moved from dark to bright.
Figure 18: Mean and standard deviation of the average absolute PCA loadings separated by layer (x-axis) for different data sets (columns) of a seven-layer BNN. Average explained variance by the first three principal components in brackets.
The traceplots displayed in Figures 19 and 20 are both examples of the fact that symmetry-induced multimodality exists and also illustrate how the within-chain variance increases in deeper layers. Moreover, Figure 20 shows how for the same setting some (in this case two) chains are “dying”, i.e., have a near-constant trace plot, while others explore the space in a meaningful way (implied by the good performance of the induced BNN).